### v1

@@ Abstract
Context. Robust impact monitoring of near-Earth objects is an essential task of planetary defense. Current systems such as NASA’s Sentry-II, University of Pisa’s CLOMON2 and Aegis of ESA have been highly successful, but independent approaches are essential to ensure reliability and to cross-validate predictions of possible impacts, probability and path on Earth.
@@
Methods. As other monitoring system NEOForCE samples several thousands of so-called ’virtual asteroids’ from the uncertainty region and integrates their orbits up to 100 years into the future. Instead of searching for close approaches of the virtual asteroids themselves with the Earth, our system looks for times when the Earth comes close to the ”realistic” uncertainty regions around them, which are mostly stretched along their osculating orbits. For each virtual asteroid we also estimate the maximal impact probability, and only if this value is large enough do we continue to the next step. In this second step, we compute how the original asteroid orbit should be slightly modified so that the new trajectory leads to an Earth impact, which allows us to confirm the possible collision and estimate the impact probability.
Results. We tested NEOForCE against NASA’s Sentry-II system on five representative asteroids with high impact probability and significant number of possible collisions: 2000 SG344, 2005 QK76, 2008 JL3, 2023 DO, and 2025 JU. NEOForCE successfully recovered mostly all possible collisions reported by Sentry-II with impact probabilities above $10^{-7}$, demonstrating the robustness of our approach. In addition, NEOForCE identified several potential impacts at the $10^{-7}$$10^{-6}$ level that Sentry-II did not report.
@@ 1 Introduction
Estimating the probability that a particular asteroid can hit the Earth in a foreseen future is an important task of planetary defense. The orbit of an asteroid in general does not have absolute or infinite precision, and for a newly discovered object the uncertainty of its orbital parameters can be significant. Moreover, this uncertainty increases with time if no additional observations are provided. Therefore, apart from the nominal solution of the orbital parameters (usually computed from the least-squares method), there are other solutions, which also well represent observations, sometimes with similar probability. These sets of orbital parameters are in general called virtual asteroids (VAs) (Milani et al., 2000). Some of these VAs can collide with the Earth and some pass by safely, which creates an impact probability different from zero or equal to one.
@@
One of the most natural and general way to estimate the impact probability is the Monte-Carlo approach: sample a large number of VAs from the uncertainty region of orbital parameters according to the probability distribution, propagate their orbit in the future, and last derive impact probability by the fraction of colliding orbits over all all considered in the sample. This approach has many advantages: It does not have additional assumptions, and also has a direct formula for estimating the uncertainty of the obtained result: $\sigma_{MC}=\sqrt{\frac{P_{MC}(1-P_{MC})}{n}}$, where $P_{MC}$ is the impact probability and $n$ is the number of considered VAs (Vavilov & Medvedev, 2015). However, if the impact probability is small—which is generally the case with PHAs—the number of required VAs to integrate to roughly estimate the impact probability (so that $\sigma_{MC}\lesssim 0.3P_{MC}$) is $10/P_{MC}$. Therefore, this approach is impractical to systematic search for possible impacts with small impact probability values on a large set of of objects.
@@
A much more effective approach is a linear approach: where one assumes that the errors of the orbital parameters at the time of a possible collision, $t$, are linearly related to the errors at the epoch of observations, $t_{0}$. In this case, one can integrate only the orbit of the nominal asteroid state vector together with variational equations, deriving the transition matrix and partial derivative matrix of orbital parameters at time $t$ over $t_{0}$. Using this matrix one can estimate the uncertainty region of the asteroid at time $t$ and assess the impact probability. The first linear method was introduced by Chodas (1993) who also proposed to use the target plane for the impact analysis. The important drawback of this approach is that the asteroid must come close to the Earth (enter its sphere of influence $\approx 929,000$ km), making linearization validity limited. Vavilov & Medvedev (2015) proposed the use of a special curvilinear coordinate system related to the asteroid’s osculating orbit to overcome this limitation. This allows us to calculate satisfactorily accurate probabilities of an asteroid collision with the Earth, even if the distance between the Earth and the nominal asteroid at time $t$ exceeds 1 au. Later Vavilov (2020) improved this approach and called it Partial Banana Mapping (PBM) method. Because the method constructs the curvilinear uncertainty region of an asteroid then finds the part of this region, which is closest to the Earth and maps it to the target plane. However, the linear relation between the errors of orbital parameters is a very important assumption and is limiting the usage of this method to the situation when the two-body formalism is not significantly violated.
@@
A better compromise between linear methods and a full Monte-Carlo approach was introduced by Milani (1999) as a concept of line of variations (LOV). Instead of sampling the full multi-dimensional phase-space, the LOV reduces the problem to a single dimension that follows the direction where uncertainties and nonlinear effects are strongest (Milani et al., 2002, 2005a, 2005b). By distributing VAs along this line and interpolating between them, one can efficiently identify possible impacting solutions and trace distinct dynamical pathways. This approach makes it possible to detect even very small impact probabilities while keeping the computational effort manageable. Because of these advantages, operational monitoring systems at NASA, Sentry [^1] https://cneos.jpl.nasa.gov/sentry/, and CLOMON2222https://newton.spacedys.com/neodys2/index.php?pc=4.1 (operated by the University of Pisa) were both built upon the LOV technique. Recently, NASA introduced the Sentry-II system (Roa et al., 2021), which samples the full confidence region rather than restricting to LOV. Potential impacts are then confirmed by adding an impact pseudo-observation and refitting the orbit to match the data. These methods require several thousands of VAs’ orbits to integrate for the first search of close encounters only, but can detect possible impacts with probabilities of the order of $10^{-7}$ while being much more efficient than the Monte-Carlo one.
@@
Vavilov & Hestroffer (2025) introduced a semi-linear method for impact probability estimation, as an extension of the linear Partial Banana Mapping method (Vavilov, 2020) built in previous work. This extension (called Impactor PBM or Improved PBM: IPBM) consists of a direct search for a state vector at the epoch of observations, $t_{0}$, that leads to a close encounter at the time of a possible collision, $t$. This method requires several orbits to be integrated, but still much less than required by Sentry of CLOMON2 systems. The method has a moderate limitation in the way that the asteroid position’s uncertainty should not exceed the whole orbit: if a VA that collides with the Earth makes different number of revolutions around the Sun than the nominal one, this collision can not be found. This naturally puts a threshold for the impact probability that can be more or less guaranteed to be found (order of $10^{-5}$). This method successfully fills the gap between restricted but very efficient linear methods and robust non-linear methods that require several thousands of orbit propagation.
@@ 1.1 Concept of the Partial Banana Mapping method
We briefly outline the concept of the Partial Banana Mapping (PBM) method, which forms the core of our system. A more detailed discussion can be found in Vavilov & Medvedev (2015); Vavilov (2020)
@@
To describe this ”banana-shaped” uncertainty region more accurately, we use a special coordinate system based on the osculating elliptical elements. In this system, one of the coordinates is the mean anomaly, $M$, and the other two $(\xi,\eta)$ describe small shifts to the side and above/below the orbit (Vavilov & Medvedev, 2015). This makes the uncertainty region easier to describe mathematically. In these curvilinear coordinates, the shape of the uncertainty region is often close to a six-dimensional Gaussian cloud.
@@ 2 NEOForCE monitoring system
Our impact monitoring system Near-Earth Object’s Forecast of Collisional Events uses a state vector (coordinates and velocities) and its 6-dimensional covariance matrix of an asteroid at the average time of observations. The covariance matrix $\mathbf{C}_{0}$ gives us an approximation of the uncertainty of the state vector $\mathbf{x}_{0}$ at time $t_{0}$. If inside this uncertainty region there is a vector of orbital parameters that leads to a collision with the Earth, then the asteroid could potentially have an impact. The goal of a monitoring system is to find all significantly large areas of orbital parameters that lead to collisions.
@@ 2.1 Initial sampling
We sample $n$ virtual asteroids (VAs) (in general $n=5001$) on this line according to Laplace distribution so that the Laplace integral ($\frac{\sqrt{2}}{2}\exp(-\sqrt{2}|x|)$) between two subsequent points is constant (and equals the integral from the last point to infinity). We use $\sqrt{2}$ to make the dispersion of the distribution equal one. Since we consider 5001 VAs the integral between the points equals $b=1/5002$. The VAs are then determined by the following:
$l_{i}=-\log(1-2ib)/\sqrt{2}*\sqrt{\lambda_{11}};\ i=-2500,2500,$ (2)
@@
We consider each VA as a representative of its own vicinity. We assign the covariance matrix to each VA by exchanging the largest eigenvalue by $\lambda_{11}^{*}=\max(l_{i}-l_{i-1},l_{i+1}-l_{i})$. The covariance matrix determining the uncertainty region for a VA is computed by:
@@ 2.2 Initial check
Instead of running NEOForCE system for all near-Earth asteroids we first fulfill the initial check whether this asteroid possibly can come close to the Earth. We select 17 VAs along LOV (on 5-sigmasigma interval equidistantly), integrate their orbits into the future and compute minimal orbit intersection distance (MOID) between their orbits and the Earth’s orbit every 15 days. If the minimal MOID among all of them is greater than a certain threshold (0.002 au), we don’t run the computation of the impact probability at all.
@@ 2.3 Close encounter precomputation
On this step for each VA we compute the information about possible close encounters. We integrate the orbit of a VA with variational equations towards the future until the final time (in general, 100 years). On each integration step we compute the osculating orbit of a VA and then obtain the minimal distance between the position of the Earth and this osculating ellipse (or parabola/hyperbola) numerically. We then find the local minima in these distances indicating the time of the local minima $t$, the state vector of the VA, $\mathbf{x}^{*}_{t}$, and the Earth and the partial derivative matrix of VA’s state vector at $t$ over original time $t_{0}$, $\mathbf{\Phi}(t_{0},t)$. The covariance matrix of the VA for time $t$ is:
@@
Then we hypothetically put the Earth at the center of the position of the VA at time $t$ (point A in Fig. 1) and compute the probability by the target plane method (Chodas, 1993). To account for the distance from the VA to the nominal orbit we then multiply this value by $\exp(-\frac{1}{2}\frac{\delta^{2}}{\lambda_{11}})$, where $\delta$ is the deviation in terms of LOV of this VA from zero (i.e. the distance on LOV of the considered VA from the nominal asteroid). This probability tells us what is the maximal possible value of the impact probability with the Earth selected asteroid can have at approximately time $t$ around the vicinity of this VA. If it is lower than our threshold (nominally $5\times 10^{-8}$) we don’t need to go to next stage to find virtual impactors. We consider this probability too low.
@@ 2.5 Search for virtual impactors
Again, we follow the idea of Milani et al. (2005a) in CLOMON2 system, that in order to confirm a possible collision, one must explicitly find a virtual impactor (a state vector at initial epoch, $t_{0}$, which actually leads to a collision). Therefore, in this stage we try to find the virtual impactor (VI) around each representative found in the previous step. For this we use a semi-linear approach extension of the PBM, briefly described here, for details see Vavilov & Hestroffer (2025).
@@
$\bm{\xi}_{*}-\bm{\xi}=\left[\frac{\partial\xi}{\partial x_{0}}\right](\bm{x}_{0*}-\bm{x}_{0}),$ (6)
@@
$(\bm{x}_{0*}-\bm{x}_{0})$ is the vector we are looking for: how to change the state vector of the selected VA on the previous stage so that after orbit integration it goes to point B at time $t$. In practice, for NEOForCE system we also consider: $\frac{3}{4}(\bm{x}_{0*}-\bm{x}_{0})$, $\frac{1}{2}(\bm{x}_{0*}-\bm{x}_{0})$ and state vectors. We integrate these orbits until time $t$ plus 0.5 days. If none of them collide with the Earth at approximately time $t$ we select the closest to the Earth. This newly selected VA is closer to the Earth at approximately time $t$ than the original one, so we repeat the procedure using this VA as the new reference (starting from computing $\bm{\xi}$, $\bm{\xi}_{*}$ vectors and matrix $\left[\frac{\partial\xi}{\partial x_{0}}\right]$). In this way, the method works as an iterative refinement process: at each step the candidate VA is moved closer to the Earth at $\approx t$ until one of three conditions is met: a collision is found, the solution drifts outside the 9-$\sigma$ uncertainty ellipsoid, or the maximum number of 10 iterations is reached.
In an extreme case where the stretching of the uncertainty region is very high and mismatches the linear estimation, the closest VA between these 3 cases can be further away than the original one. In this case we select $\frac{1}{6}(\bm{x}_{0*}-\bm{x}_{0})$ virtual asteroid and proceed again. Once the VA entered the sphere of influence (approximately 929,000 km) we launch the target plane analysis (see Section 3).
@@
Figure 3: Schematic illustration of the work of NEOForCE system. On the upper chart one can see the nominal asteroid (purple dot), the original uncertainty region at the epoch of observation (black ellipse), the LOV (horizontal line), VAs sampled along LOV (black dots). The green area represents the smaller uncertainty region assigned to this VA. The red denotes the VI. On the bottom chart one can see the nominal position of the asteroid at time of a possible collision, $t$ (point A), the osculating orbit of the nominal asteroid (black part of an ellipse), the Earth, the position of the VA at time $t$ (green dot), the uncertainty region assigned for this VA at time $t$ (green curved ellipse). The red dot (point B) represents the position of the VI. The dashed vector shows the relative motion of the Earth and the green uncertainty region. The Earth collides with the red dot.
A schematic illustration of the work of our system is presented in Fig. 3.

### v2

@@ Abstract
Context. Robust impact monitoring of near-Earth objects is an essential task of planetary defense. Current systems such as NASA’s Sentry-II, NEODyS’s CLOMON2 and Aegis of ESA have been highly successful, but independent approaches are essential to ensure reliability and to cross-validate predictions of possible impacts, probability and path on Earth.
@@
Methods. As other monitoring system NEOForCE samples several thousands of so-called ’virtual asteroids’ from the uncertainty region and integrates their orbits up to 100 years into the future. Instead of searching for close approaches of the virtual asteroids themselves with the Earth, our system looks for times when the Earth comes close to the ”realistic” uncertainty regions around them, which are mostly stretched along the osculating ellipses of virtual asteroids. For every virtual asteroid and every possible collision time we also estimate the maximal impact probability, and only if this value is large enough ($5\times 10^{-8})$ we continue to the next step. In this second step, we compute how the original asteroid orbit should be slightly modified so that the new trajectory leads to an Earth impact, which allows us to confirm the possible collision and estimate the impact probability.
Results. We tested NEOForCE against NASA’s Sentry-II system on five representative asteroids with high impact probability and significant number of possible collisions: 2000 SG344, 2005 QK76, 2008 JL3, 2023 DO, and 2008 EX5. NEOForCE successfully recovered mostly all possible collisions reported by Sentry-II with impact probabilities above $10^{-7}$, demonstrating the robustness of our approach. In addition, NEOForCE identified several potential impacts at the $10^{-7}$$10^{-6}$ level that Sentry-II did not report.
@@ 1 Introduction
Estimating the probability that a particular asteroid can hit the Earth in a foreseen future is an important task of planetary defense. The orbit of an asteroid in general does not have absolute or infinite precision, and for a newly discovered object the uncertainty of its orbital parameters can be significant. Moreover, this uncertainty increases with time if no additional observations are provided. Therefore, apart from the nominal solution of the orbital parameters (usually computed from the least-squares method), there are other solutions, which also well represent observations, sometimes with similar probability. These sets of orbital parameters are in general called virtual asteroids (VAs) (Milani et al., 2000). Some of these VAs can collide with the Earth and some pass by safely, which allows the impact probability to be in a range between 0 and 1.
@@
One of the most natural and general way to estimate the impact probability is the Monte-Carlo approach: sample a large number of VAs from the uncertainty region of orbital parameters according to the probability distribution, propagate their orbit in the future, and last derive impact probability by the fraction of colliding orbits over all considered in the sample. This approach has many advantages: It does not have additional assumptions, and also has a direct formula for estimating the uncertainty of the obtained result: $\sigma_{MC}=\sqrt{\frac{P_{MC}(1-P_{MC})}{n}}$, where $P_{MC}$ is the impact probability and $n$ is the number of considered VAs (Vavilov and Medvedev, 2015). However, if the impact probability is small—which is generally the case with NEAs —the number of required VAs to integrate to roughly estimate the impact probability (so that $\sigma_{MC}\lesssim 0.3P_{MC}$) is $10/P_{MC}$. Therefore, this approach is impractical to systematic search for possible impacts with small impact probability values on a large set of of objects.
@@
A much more effective approach is a linear approach: where one assumes that the errors of the orbital parameters at the time of a possible collision, $t$, are linearly related to the errors at the epoch of observations, $t_{0}$. In this case, one can integrate only the orbit of the nominal asteroid state vector together with variational equations, deriving the transition matrix and partial derivative matrix of orbital parameters at time $t$ over $t_{0}$. Using this matrix one can estimate the uncertainty region of the asteroid at time $t$ and assess the impact probability. The first linear method was introduced by Chodas (1993) who also proposed to use the target plane for the impact analysis. In addition to the linear assumption (which can be broken by a close encounter) the important drawback of this approach is that the asteroid must come close to the Earth (enter its sphere of influence $\approx 929,000$ km), making linearization validity limited. Vavilov and Medvedev (2015) proposed the use of a special curvilinear coordinate system related to the asteroid’s osculating orbit to overcome this limitation. This allows us to calculate satisfactorily accurate probabilities of an asteroid collision with the Earth, even if the distance between the Earth and the nominal asteroid at time $t$ exceeds 1 au. Later Vavilov (2020) improved this approach and called it Partial Banana Mapping (PBM) method. Because the method constructs the curvilinear uncertainty region of an asteroid then finds the part of this region, which is closest to the Earth and maps it to the target plane. However, the linear relation between the errors of orbital parameters is a very important assumption and is limiting the usage of linear methods to the situation when the two-body formalism is not significantly violated.
@@
A better compromise between linear methods and a full Monte-Carlo approach was introduced by Milani (1999) as a concept of line of variations (LOV). Instead of sampling the full multi-dimensional phase-space, the LOV reduces the problem to a single dimension that follows the direction where uncertainties and nonlinear effects are strongest (Milani et al., 2002, 2005b, 2005a). By distributing VAs along this line and interpolating between them, one can efficiently identify possible impacting solutions and trace distinct dynamical pathways. This approach makes it possible to detect even very small impact probabilities while keeping the computational effort manageable. Because of these advantages, operational monitoring systems at NASA, Sentry [^1] https://cneos.jpl.nasa.gov/sentry/, CLOMON2222https://newton.spacedys.com/neodys2/index.php?pc=4.1 (operated by NEODyS) and ESA’s Aegis [^3] https://neo.ssa.esa.int/risk-list were all built upon the LOV technique. Recently, NASA introduced the Sentry-II system (Roa et al., 2021), which samples the full confidence region rather than restricting to LOV. Potential impacts are then confirmed by adding an impact pseudo-observation and refitting the orbit to match the data. These methods require several thousands of VAs’ orbits to integrate for the first search of close encounters only, but can detect possible impacts with probabilities of the order of $10^{-7}$ while being much more efficient than the Monte-Carlo one.
@@
Vavilov and Hestroffer (2025) introduced a semi-linear method for impact probability estimation, as an extension of the linear Partial Banana Mapping method (Vavilov, 2020) built in previous work. This extension (called Impactor PBM or Improved PBM: IPBM) consists of a direct search for a state vector at the epoch of observations, $t_{0}$, that leads to a close encounter at the time of a possible collision, $t$. This method requires several orbits to be integrated, but still much less than required by Sentry of CLOMON2 systems. The method has a moderate limitation in the way that the asteroid position’s uncertainty should not exceed the whole orbit: if a VA that collides with the Earth makes different number of revolutions around the Sun than the nominal one, this collision can not be found. This naturally puts a threshold for the impact probability that can be more or less guaranteed to be found (order of $10^{-5}$). This method successfully fills the gap between restricted but very efficient linear methods and robust non-linear methods that require several thousands of orbit propagation.
@@ 1.1 Concept of the Partial Banana Mapping method
We briefly outline the concept of the Partial Banana Mapping (PBM) method, which forms the core of our system. A more detailed discussion can be found in Vavilov and Medvedev (2015); Vavilov (2020).
@@
To describe this ”banana-shaped” uncertainty region more accurately, we use a special coordinate system based on the osculating elliptical elements. In this system, one of the coordinates is the mean anomaly, $M$, and the other two $(\xi,\eta)$ describe small shifts to the side and above/below the orbit (Vavilov and Medvedev, 2015). This makes the uncertainty region easier to describe mathematically. In these curvilinear coordinates, the shape of the uncertainty region is often close to a six-dimensional Gaussian cloud.
@@ 2 NEOForCE monitoring system
Our impact monitoring system Near-Earth Object’s Forecast of Collisional Events uses a state vector (coordinates and velocities) and its $6\times 6$ covariance matrix of an asteroid at the average time of observations. The covariance matrix $\mathbf{C}_{0}$ gives us an approximation of the uncertainty of the state vector $\mathbf{x}_{0}$ at time $t_{0}$. If inside this uncertainty region there is a possible initial state vector that leads to a collision with the Earth, then the asteroid could potentially have an impact. The goal of a monitoring system is to find all significantly large areas of state vectors that lead to collisions, later referred as Virtual Impactors (VI).
@@ 2.1 Initial sampling
We sample $n$ VAs (in general $n=5001$) on this line according to Laplace distribution so that the Laplace integral ($\frac{\sqrt{2}}{2}\exp(-\sqrt{2}|x|)$) between two subsequent points is constant (and equals the integral from the last point to infinity). We use $\sqrt{2}$ to make the dispersion of the distribution equal one. Since we consider $n$ VAs the integral between the points equals $b=1/(n+1)$. The VAs are then determined by the following:
$l_{i}=-\frac{1}{\sqrt{2}}\log(1-2ib)\sqrt{\lambda_{11}};\ \ l_{-i}=-l_{i};\ \ i=0,{n/2}$ (2)
@@
We consider each VA as a representative of its own vicinity. We assign the covariance matrix to each VA by exchanging the largest eigenvalue by $\lambda_{11}^{*}=\max(l_{i}-l_{i-1},l_{i+1}-l_{i})$. Thus, a 1-sigma uncertainty ellipsoid around each VA would slightly touch or even include the neighbor VA. The covariance matrix determining the uncertainty region for a VA is computed by:
@@ 2.2 Initial check
Figure 2: Upper chart. The histogram of minimal MOID value for asteroids presented on CNEOS web-site as ones that have non-zero impact probabilities with the Earth on September 9, 2025. Bottom chart. The cumulative histogram of minimal MOID for all NEOs. The vertical red line is our empirical 0.02 au threshold.
Instead of running NEOForCE system for all near-Earth asteroids we first fulfill the initial check whether this asteroid possibly can come close to the Earth. We select 100 VAs along LOV (on 5-$\sigma$ interval equidistantly), integrate their orbits into the future and compute minimal orbit intersection distance (MOID) between their orbits and the Earth’s orbit every 100 days. If the minimal MOID among all of them is greater than a certain threshold (0.02 au), we don’t run the computation of the impact probability at all. Fig. 2 presents this minimal MOID value for all the asteroids, which are in the impact risk table on CNEOS web-site (on September 9, 2025), as well as the minimal MOID for all NEOs. As one can see our threshold 0.02 au includes all currently known NEAs with non-zero impact probability. At the same time it allows us to limit the number of launches of the monitoring system to approximately 15,000 instead of 40,000. This initial check requires less than one minute of computations for one asteroid on one thread of a processor like Intel(R) Xeon(R) Gold 6230 2.10GHz.
@@ 2.3 Close encounter precomputation
On this step for each VA we compute the information about possible close encounters. We integrate the orbit of a VA with variational equations towards the future until the final time (in general, 100 years). On each integration step we compute the osculating orbit of a VA and then obtain the minimal distance between the position of the Earth at this time and the points of this osculating ellipse (or parabola/hyperbola) numerically. We then find the local minima in these distances indicating the time of the local minima $t$, the state vector of the VA, $\mathbf{x}^{*}_{t}$, and the Earth and the partial derivative matrix of VA’s state vector at $t$ over original time $t_{0}$, $\mathbf{\Phi}(t_{0},t)$. The covariance matrix of the VA for time $t$ is:
@@
Then we hypothetically put the Earth at the center of the position of the VA at time $t$ (point A in Fig. 1) and compute the probability by the target plane method (Chodas, 1993). To account for the distance from the VA to the nominal orbit we then multiply this value by $\exp(-\frac{1}{2}\frac{l_{i}^{2}}{\lambda_{11}})$, where $l_{i}$ is from Eq. 2. This probability tells us what is the maximal possible value of the impact probability with the Earth selected asteroid can have at approximately time $t$ around the vicinity of this VA. If it is lower than our threshold (nominally $5\times 10^{-8}$) we don’t need to go to next stage to find virtual impactors. We consider this probability too low.
@@ 2.5 Search for virtual impactors
Again, we follow the idea of Milani et al. (2005b) in CLOMON2 system, that in order to confirm a possible collision, one must explicitly find a virtual impactor (a state vector at initial epoch, $t_{0}$, which actually leads to a collision). Therefore, in this stage we try to find the virtual impactor (VI) around each representative found in the previous step. For this we use a semi-linear approach extension of the PBM, briefly described here, for details see Vavilov and Hestroffer (2025).
@@
$\bm{\xi}_{*}-\bm{\xi}=\left[\frac{\partial\bm{\xi}}{\partial\bm{x_{0}}}\right](\bm{x}_{0*}-\bm{x}_{0}),$ (6)
@@
$(\bm{x}_{0*}-\bm{x}_{0})$ is the vector we are looking for: how to change the state vector of the selected VA on the previous stage so that after orbit integration it goes to point B at time $t$. In practice, for NEOForCE system we also use these two additional state vectors: $\frac{3}{4}(\bm{x}_{0*}-\bm{x}_{0})+\bm{x}_{0}$ and $\frac{1}{2}(\bm{x}_{0*}-\bm{x}_{0})+\bm{x}_{0}$. These state vectors are closer to $\bm{x}_{0}$ than $\bm{x}_{0*}$. We integrate these orbits until time $t$ plus 0.5 days. If none of them collide with the Earth at approximately time $t$ we select the closest to the Earth. This newly selected VA is closer to the Earth at approximately time $t$ than the original one, so we repeat the procedure using this VA as the new reference (starting from computing $\bm{\xi}$, $\bm{\xi}_{*}$ vectors and matrix $\left[\frac{\partial\xi}{\partial x_{0}}\right]$). In this way, the method works as an iterative refinement process: at each step the candidate VA is moved closer to the Earth at $\approx t$ until one of four conditions is met:
  1. (i)
    a collision is found: we compute the probability by the target plane method (see Section 3) ;
  2. (ii)
    the solution drifts outside the 9-$\sigma$ uncertainty ellipsoid: we stop the procedure
  3. (iii)
    the maximum number of 10 iterations is reached: we stop the procedure;
  4. (iv)
    the asteroid enters the sphere of influence with the Earth — we run the target plane analysis (see Section 3) to search for a VI.
By taking two additional state vectors, which are closer to original one, instead of simply using $\bm{x}_{0*}$ we slightly mitigate some convergence problems that can happen after multiple close encounters. In an extreme case where the stretching of the uncertainty region is very high and mismatches the linear estimation, the closest VA between these 3 cases can be further away from the Earth than the original one. In this case we select $\frac{1}{6}(\bm{x}_{0*}-\bm{x}_{0})+\bm{x}_{0}$ VA and proceed again.
@@
Figure 4: Schematic illustration of the work of NEOForCE system. On the upper chart one can see the nominal asteroid (purple dot), the original uncertainty region at the epoch of observation (black ellipse), the LOV (horizontal line), VAs sampled along LOV (black dots). The green area represents the smaller uncertainty region assigned to this VA. The red denotes the VI. On the bottom chart one can see the nominal position of the asteroid at time slightly before a possible collision, $t$ (point A), the osculating orbit of the nominal asteroid (black part of an ellipse), the Earth, the position of the VA (green dot), the uncertainty region assigned for this VA at time slightly before $t$ (green curved ellipse). The red dot (point B) represents the position of the VI. The dashed vector shows the relative motion of the Earth and the green uncertainty region. The Earth collides with the red dot.